- Site is muddy due to ongoing rain
- Smoke is less visible than yesterday
- 40 loads of material from the borrow pit have been moved to the site.
- Grading on the southwest slope and more work on the northeast area of the site is continuing today.
- More regulated solid waste was observed in the staging area.